OpenStudy (anonymous):

Geometry word problem help! triangle ABC is similar to triangle FGH with medians AP and FQ to sides BC and GH respectively, AB=9, and FG=6. If AP is 4 greater than FQ, find both medians.

OpenStudy (rational):

draw it out

OpenStudy (anonymous):



OpenStudy (rational):

Hint : medians also scale by same factor when triangles are similar

OpenStudy (rational):

\[\large \dfrac{9}{6} = \dfrac{AP}{FQ}\]

OpenStudy (anonymous):

Hmmm. I see. Thanks!

OpenStudy (anonymous):

But AP is 4 greater than FQ?

OpenStudy (rational):

Yes so \(AP = FQ+4\)

OpenStudy (rational):

\[\large \dfrac{9}{6} = \dfrac{FQ+4}{FQ}\] cross multiply and solve \(FQ\) first

OpenStudy (anonymous):

FQ=8?

OpenStudy (rational):

Yep

OpenStudy (anonymous):

so AP =32?

OpenStudy (rational):

how ? AP = FQ + 4 = 8 + 4 = 12 right

OpenStudy (anonymous):

so AP = 12

OpenStudy (anonymous):

I see. I multiplied instead.